OpenStudy (anonymous):

Solve: 3c + 2 = 4c - 9.

OpenStudy (anonymous):

3c+2=4c-9 3c-4c=-9-2 -c=-11 c=11

OpenStudy (anonymous):

Just algebraically manipulate the equation to separate the variable. Then combine like terms. and find the solution for c c=11